Hercules Exercise Equipment Co. purchased a computerized measuring device two years ago for $90,000. The equipment falls into the five-year category for MACRS depreciation and can currently be sold for $40,800. A new piece of equipment will cost $290,000. It also falls into the five-year category for MACRS depreciation. Assume the new equipment would provide the following stream of added cost savings for the next six years. Use Table 12-12. Use Appendix B for an approximate answer but calculate your final answer using the formula and financial calculator methods. Cash Ycar savings 1 S67,000 57,000 55,000 53,000 50,000 39,000 6 The firms tax rate Is 40 percent and the cost of capltal Is 14 percent. a. What is the book value of the old equipment? (Do not round intermediate calculations and round your answer to the nearest whole dollar.) Book value b. What is the tax loss on the sale of the old equipment? (Do not round intermediate calculations and round your answer to the nearest whole dollar.) Tax loss c. What is the tax benefit from the sale? (Do not round intermediate calculations and round your answer to the nearest whole dollar.) Tax benefi

d. What is the cash inflow from the sale of the old equipment? (Do not round intermediate calculations and round your answer to the nearest whole dollar.) Cash intlow e. What is the net cost of the new equipment? (Include the inflow from the sale of the old equipment.) (Do not round intermediate calculations and round your answer to the nearest whole dollar.) Net cost f. Determine the depreciation schedule for the new equipment. (Round the depreciation base and annual depreciation answers to the nearest whole dollar. Round the percentage depreciation factors to 3 decimal places.) Depreciation Percentage Annual Base Depreciation Depreciation

g. Determine the depreciation schedule for the remaining years of the old equipment.(Round the depreciation base and annual depreciation answers to the nearest whole dollar. Round the percentage depreciation factors to 3 decimal places.) PercentageDepr Depreciation Annual Base eciation h. Determine the incremental depreciation between the old and new equipment and the related tax shield benefits. (Enter the tax rate as a decimal rounded to 2 decimal places. Round all other answers to the nearest whole dollar.) Depreciation Depreciation Incremental on Old Equipment Depreciation Tax Rate Tax Shleld Benefits Year on New

i. Compute the aftertax benefits of the cost savings. (Enter the aftertax factor as a decimal rounded to 2 decimal places. Round all other answers to the nearest whole dollar.) Aftertax 1- Tax Rate) Savings Year Savings 67.000 57,000 55,000 53,000 50000 39000 j-1. Add the depreciation tax shield benefits and the aftertax cost savings to determine the total annual benefits. (Do not round intermediate calculations and round your answers to the nearest whole dollar. Tax Shield Aftertax Cost Total Annua Year Benefits from Savings Depreciation

j-2. Compute the present value of the total annual benefits. (Do not round intermediate calculations and round your answer to the nearest whole dollar.) Total annual benents k-1. Compare the present value of the incremental benefits ω to the net cost of the new equipment(e), (Do not round intermediate calculations. Negative amount should be indicated by a minus sign. Round your answer to the nearest whole dollar.) Total annual benefits k-2. Should the replacement be undertaken? O Yes
Answer #1

a. Purchase price = 90,000

MACRS 5 year depreciation for first two year = 20% = 32% = 52%

Book Value = 90,000*(1-0.52) = 43,200

Book value of the old Machine = $43,200

b. Tax loss on Sale = Book Value - Sale price =43,200 -40,800 = 2,400

Tax loss on Sale =$2,400

c. Tax benefit from the sale = Loss * tax rate = 2,400 * 0.4 = 960

Tax benefit from the sale = $960

d. Cash inflow from sale of old equipment = 40,800 + 960 = 41,760

Cash inflow from sale of old equipment = 41,760

Note: We have answered four sub-parts. Only four sub-parts can be answered at a time. Please post others separately for experts to answer